OpenStudy (anonymous):

Rationalize the denominator.. \[\frac{ -12x }{ \sqrt{x-10} }\] @kelliegirl33

OpenStudy (anonymous):

@eliassaab @hba @hartnn @geneshie8 @robtobey @thomaster @Opcode

OpenStudy (anonymous):

\[\frac{ -12x \times \sqrt{x-10} }{ x-10}\]

OpenStudy (anonymous):

So does that cancel out \[x-10\] All together?

OpenStudy (anonymous):

nope, i will think abt it, gtg, will definitely give u an answer after i went back from lunch

OpenStudy (anonymous):

Okay, thanks mate :)

OpenStudy (anonymous):

@Loser66 @AravindG @AccidentalAiChan @kelliegirl33

OpenStudy (anonymous):

Can anyone help please? =/.. Been waiting for over an hour now..

OpenStudy (loser66):

I saw you said "Thanks mate" doesn't it mean you got it?

OpenStudy (anonymous):

@trollwillcontinue stated "nope, i will think abt it, gtg, will definitely give u an answer after i went back from lunch" And considering he said he ".. will definitely give u an answer after I went back from lunch" I used manors and stated "Thanks mate".

OpenStudy (anonymous):

Considering he was offering to help.

OpenStudy (loser66):

his first comment was the correct answer, I can do no more for that perfect answer

OpenStudy (anonymous):

Oh, that's the answer? Lol didn't know. Sorry :P

OpenStudy (loser66):

hihihhi

OpenStudy (anonymous):

i tried to remove the unknown but i realised all methods do not work and hence the final rationalised number should be the first one. sorry for taking so long. gtg sleeping

Directrix (directrix):

@trollwillcontinue Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution http://openstudy.com/code-of-conduct